an archer fires a 0.30-kg arrow from a bow .the archer exerts an average force of 300 N to draw the string back 0.50 m


1
Expert's answer
2020-10-29T07:02:48-0400
"W=Fd=0.5mv^2\\(300)(0.5)=0.5(0.3)v^2\\v=32\frac{m}{s}"